OpenStudy (shes.radical):

Choose the point-slope form of the equation below that represents the line that passes through the point (6, -3) and has a slope of 1/2. A) Y-6=1/2(x+3) B) Y=1/2x-6 C) Y+3=1/2(x-6) D) X-2Y=12

OpenStudy (aripotta):

point-slope form is y - y1 = m(x - x1), where m is slope. just plug your info into the formula to get your answer

OpenStudy (shes.radical):

I dont get it../)-(\

OpenStudy (shes.radical):

Wait is it A??

OpenStudy (aripotta):

the problem gives us the point (6,-3). points are written in the form (x,y) so our y1 will be -3, and our x1 will be 6

OpenStudy (aripotta):

and it also gives us the slope as 1/2. once we plug that info in, we get y + 3 = 1/2(x - 6)

OpenStudy (shes.radical):

So its C..ahh you rock!! Thanks!

OpenStudy (aripotta):

\(\Large \boldsymbol{ \rlap{\color{#008AFF}{\hspace{2.5pt}n}\color{#0073FF}{\hspace{0.34pt}o} \hspace{3pt}\color{#005CFF}{p} \color{#0045FF}{\hspace{0.9pt}r} \color{#0017FF}{o} \color{#0200FF}{\hspace{1.7pt}b} \color{#3000FF}{\hspace{1.5pt}l} \color{#4700FF}{\hspace{0.4pt}e} \color{#7500FF}{\hspace{0.9pt}m} \color{#8C00FF}{\hspace{0.5pt}o}~\LARGE\color{#D100FF}{\ddot\smile}}{\hspace{1.5pt} no~p\hspace{0.3pt}r\hspace{0.4pt}o\hspace{1pt}b\hspace{1.5pt}l\hspace{0.4pt}e\hspace{0.4pt}m\hspace{0.4pt}o}}\) don't forget to give me best response :P lol